the equivalent weight of metal is 30 what volume of chlorine would combine at 27 degree Celsius and 750 M pressure with 0.6 gram of metal?(aquous tention at 27°C = 26.6mm

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
In this response , mole likeness metal = mole likeness Cl 
Mole eq of metal = 0.6/30 = 0.02 mol 
Same will be Cl 
So Cl is responding 0.02 mol eq = 0.01 mol ( as Cl2 when respond gives 2 mol Cl-... i.e. n factor =2) 
.01 mol Cl in 27°C , 750M press ( i.e. STP) is (0.01*22.4L)= 0.224 L .
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)